Opened 4 years ago

Closed 3 years ago

#581 closed flight processing (fixed)

EUFAR15/38, flight day 169a/2015, Mallorca

Reported by: dac Owned by:
Priority: immediate Milestone: 2015 data processing completion
Component: Processing: general Keywords:
Cc: Other processors:

Description

Data location: ~arsf/arsf_data/2015/flight_data/spain/EUFAR15_38-2015_169a_Mallorca

Data arrived from ARSF via SATA disk on 09/07/2015

Scientific objective: Use airborne hyperspectral, LiDAR and SfM to improve process understanding by monitoring and modelling changing patterns of ecogeomorphological connectivity in Mediterranean insular catchments.

Priority: Unknown

PI: Joan Estrany
EUFAR Project ID: MEDhy2CON

Any other notes..

Sensors:

  • Fenix (requested)
  • Leica LIDAR (requested)
  • Leica FW LIDAR (not requested but flown anyway)
  • Owl (not requested, only available for 4 lines)

Change History (27)

comment:1 Changed 4 years ago by dac

Couldn't generate logsheet using 'generateLogsheet.py' problem due to unassigned variable.

comment:2 Changed 4 years ago by asm

Problem with script 'generateLogsheet.py' has been solved. Logsheet has been generated now. There are 13 flightlines but the Owl failed during the journey.

comment:3 Changed 4 years ago by asm

Navigation
Starting navigation process.
Basestation data needed teqc to be processed.

Lat: 39 32 26.25388
Lon: 2 43 19.99818
Elipsoidal Height: 53.556 metres
Antenna Height: 0.085 metres

Navigation Processing
Started with IpasTC.

Last edited 4 years ago by asm (previous) (diff)

comment:4 Changed 4 years ago by asm

Navigation Processing
Finished. Everything looks OK.

comment:5 Changed 4 years ago by asm

Hyperspectral Processing
Starting fenix processing

comment:6 Changed 4 years ago by asm

Hyperspectral Processing

Could not 'generate_apl_config.py' there was a problem with flightlines number 7 and 11. The data has been moved to folder /hyperspectral/fenix/problematic_flightlines

We are not processing those two flightlines for now.

comment:7 Changed 4 years ago by asm

Hyperspectral Processing

Dac updated the boresight csv for this flight. Thank you Dac.

Have the same problem that 160 2015, error message:
Requested sync time index is out of bounds in GetSyncDelay()
needed to set on config file:
use_nav = false

Last edited 4 years ago by asm (previous) (diff)

comment:8 Changed 4 years ago by asm

Hyperspectral Processing
Flightlines 01 and 02 are apart from the other flightlines.

Found SCT values

Flightline FENIX
1 3.30
2 3.36
3 2.12
4 3.27
5 4.34
6 2.28
7 N/A
8 2.67
9 3.16
10 4.39
11 N/A
12 3.00
13 3.49
Last edited 4 years ago by asm (previous) (diff)

comment:9 Changed 4 years ago by asm

Hyperspectral Processing
Awaiting new fenix calibration

comment:10 Changed 4 years ago by gej

Started Lidar Processing

comment:11 Changed 4 years ago by gej

Lidar Processing

Roll and pitch vary through flightlines
Flightlines 1 and 2 away from rest of group.

FlightlineRollPitch
1 218-0.00095-0.0022
2 949-0.00095-0.0022
3 509-0.00110-0.00225
4 018-0.00110-0.00225
5 538-0.00090-0.0018
6 154-0.00108-0.0023
7 829-0.00100-0.0022
8 537-0.00120-0.00225
9 407-0.0006-0.00230
10 103-0.00115-0.00240
11 736-0.0006-0.00228
12 446-0.00120-0.00232
13 225-0.001-0.00236
Last edited 4 years ago by gej (previous) (diff)

comment:12 Changed 4 years ago by gej

Lidar Processing

Discrete Classification complete

Starting fullwave form

comment:13 Changed 4 years ago by gej

Lidar Processing

Full waveform processing complete,

Creating Delivery

comment:14 Changed 4 years ago by gej

Lidar Processing

Delivery creation complete, Ready for DC

comment:15 Changed 4 years ago by stgo

Lidar Delivery Check

Starting check

Readme:
"line 027 Elevation difference could not be calculated due to high amount of vegetation and mountainous regions."

Should read:
"The elevation difference of line 027 could not be calculated due to high levels of vegetation and mountainous regions."

Need to specify N/S in projection

Screenshots:
Line 01 and 02 both suffer from AGC striping (I think), it would be worth mentioning this somewhere.

Logsheet and intensity image are incorrectly named according to proj_tidy

Last edited 4 years ago by stgo (previous) (diff)

comment:16 Changed 4 years ago by gej

Lidar Processing

Have made suggested change to Elevation difference comment.
Specified N in projection.
Added small comment about line 001 and 002 suffering from AGC striping to readme

Renamed logsheet.
Have not changed intensity screenshot as after checking proj_tidy it mentions the intensity vectors screenshot, not the intensity screenshot.

comment:17 Changed 4 years ago by stgo

Lidar Delivery Check

Looks good, ready for delivery.

comment:18 Changed 4 years ago by asm

Lidar Data Delivery
Delivered on FTP server. Sent notification to PI.

comment:19 Changed 4 years ago by asm

Hyperspectral Processing
Delivery created. It still needs the updated data quality report but the project is ready for starting with Delivery Check.

P.S.: New data quality report has been added to the delivery

Last edited 4 years ago by asm (previous) (diff)

comment:20 Changed 4 years ago by lah

Started Fenix DC
Do we have a reason why lines 7 & 11 could not be processed that can be put in the readme? Also, the table on page 3 needs these 2 lines removing (Logsheet names 033 & 029).

comment:21 Changed 4 years ago by asm

Fenix Processing
Readme updated.

comment:22 Changed 4 years ago by lah

Fenix DC

  • Renamed logsheet to match lidar delivery
  • Under/overflows look ok, but lines 1,5,6 &,8 have a strip of underflows/missing data at the top of the line for all bands which isn't mentioned.(probably too few for automated script) Looking at the data logs, this is due to dropped frames between lines 1-28. Have added to readme.
  • spectra look ok. There is consistently a peak at end of VNIR for all flight lines, but this is in the masked out region.
  • Have also regenerated the new readme to get rid of "?? section" problem.

Ready to deliver once files have zipped.

comment:23 Changed 4 years ago by lah

Fenix ready to deliver, files all zipped.

comment:24 Changed 4 years ago by gej

Hyperspectral Data Delivered to PI on 20th October 2015 via FTP

comment:25 Changed 4 years ago by dac

Archiving

Started Archiving.

comment:26 Changed 4 years ago by gej

PI had problems with download data via FTP so has now been delivered via HDD to University of Balearic Islands, 18th November 2015

comment:27 Changed 3 years ago by dac

  • Resolution set to fixed
  • Status changed from new to closed

Archiving

Data now available via NEODC.

Note: See TracTickets for help on using tickets.